OpenStudy (anonymous):

Multiply the polynomials: 1. (a + 3)(a - 2) 2. (x2 + 4)(x2 - 4) 3. (x2 + 3x + 1)(x2 + x + 2)

OpenStudy (anonymous):

use the distributive property \[a(b\pm c)=ab\pm ac\]

OpenStudy (anonymous):

using this you take the first variable of the first and multiply it by the second parenthesis and same with the second variable/constant of the first \[a(a-2)+3(a-2)\]

OpenStudy (anonymous):

what do you get when yo udo this?

OpenStudy (anonymous):

So then i got: 2a-2a + 3a-6

OpenStudy (anonymous):

But I dont think that is right lol

OpenStudy (anonymous):

@Outkast3r09

OpenStudy (anonymous):

everything but your first multiplication is correct \[a(a)\neq 2a\]

OpenStudy (anonymous):

hint: when you have a number to a power you get this correct \[2^2=2*2\]

OpenStudy (anonymous):

so \[a*a=?\]

OpenStudy (anonymous):

uhhh 2 squared ???

OpenStudy (anonymous):

@Outkast3r09

OpenStudy (anonymous):

not quite so 2 is the base

OpenStudy (anonymous):

in \[2^2\]

OpenStudy (anonymous):

the power tells you how many times you multiply that base by itself

OpenStudy (anonymous):

so \[2^3=2*2*2\] \[2^5=2*2*2*2*2\]

OpenStudy (anonymous):

so in your case \[ a*a\] a is your base and the number of bases you multiply is your power

OpenStudy (anonymous):

so \[a*a=a^2\]

OpenStudy (anonymous):

I know how to do powers lol. I just dont know what \[a \times a\] is

OpenStudy (anonymous):

oh okay lol

OpenStudy (anonymous):